FEL Upgrade Project Weekly Brief
Decemer 18-22, 2000


On behalf of the Jefferson Lab FEL Team , we thank all of our stakeholders at the Dept. of the Energy, the Office of Naval Research, the Commonwealth of Virginia and the FEL user community for their support, advice ,criticism and encouragement during 2000.  We wish all of you a pleasant holiday season and a good new year .

2000  was a good year for all of us involved in the FEL program . For all of our collaborators I offer  a short synopsis:

1. We provided over 1000 hours of user time which generated world class scientific results in materials science, biophysics, nanotechnology and microfabrication.

2. The FEL provided new power records for sub-picosecond class laser light at 1 micron (>300 watts).

3. We demonstarted wiggler efficiencies for converting electron power to optical power of over 2%.

4. We generated  sub-picosecond, high brightness  x-rays in the 5 kV range using Compton scattering.

5. We launched the ONR sponsored 10 kW IR Upgrade project in June and the project is off to healthy start for delivering an upgraded machine in 2002.

6. We anticipate funding in early 2001 by the Air Force Research Lab of a 1kW UV upgrade for the FEL for materials processsing applications.

7. One of our Laser Processsing Consortium industrial partners donated a compact, superconducting
synchrotron x-ray source to SURA/Jefferson Lab which will be integrated into a planned expansion of the FEL User Facility


The project monthly report for November 2000 including the technical and cost-performance summaries were sent to the DOE contracting officer and ONR contract monitor.

DOE, ONR and Jefferson Lab completed negotiation on the MOA and statement of work for the FY2001  contract that would cover completion of the 10 kW IR Upgrade project in FY2001-2002. Copies of the MOA with DOE signatures were forwarded to ONR for signature.

We distributed the preliminary agenda for the next Laser Processsing Consortium/FEL Users Workshop which will be held on Jan.18-19.  The agenda has also been posted on the FEL web site(http://www.jlab.org/FEL/LPC/lpcfirstannounc1-01.html)

WBS 3 (Beam Physics):

Optimization/fine-tuning of Revision 1.1 design solution continued.
"Final" solution features:

1. revised injection line design using sector dipoles
2. "high probability" turn on energy gains (40 MeV, 55 MeV, 40 MeV) with associated linac optics
    solution (larger injection beam envelopes)
3. additional drift space in endloops for correction elements
4. GW parameters common to IR and UV systems

Discussions with beam transport engineering (WBS 9) led to agreement on documentation format for magnet specification.

WBS 4 (Injector):

Completed construction of a reverse flow heater for cleaning of the GaAs.  A test will be run on the current wafer in service for the FEL to determine if a more uniform quantum efficiency can be achieved with the cathode. We began looking at possible lay-out configurations in the Injector Test Stand that will allow us to test the upgraded FEL gun at the end of 2001 without intefering with current activities with the nuclear phsyics gun tests.

WBS 5 (SRF):

In the cavity area, we are continuing to fabricate cavity segments, weld them up and prepare them for cavity tuning.  A vendor visit was made to the waveguide supplier for kick off of the contract.  The vendor looks capable and well suited to supply parts.  The first article helium vessels were shipped this month from PHPK, the vessel vendor.

WBS 6 (RF):

We present a somewhat longer summary of the RF system this week in order to fully document the present status of these systems.

Zone 4 installation and testing:

Completed alignments and checks of the cathode currents and body currents for all eight positions.  We found that the new analog instrumentation boards do not have the input diodes that protect the op
amp, connected to -15 volts.  This caused the KBCU (body current) to fault anytime current goes greater than 600mA.  For FL04, the diodes have been removed.  We will need to look at the board to see where and how to connect the -15 volts for the other boards.

We had the CPS up and running with all eight klystrons connected but the mod anode boards pinched off all of the tubes.  It appears that the transformer did not have enough windings between the primary and secondary to turn on the FETs.  This problem should be resolved (thanks to Mark Augustine and Dave Gelhaar).

Unfortunately we will not be able to turn the zone on again until sometime next year.  There was a water leak somewhere in the LCW system and in the process of digging to investigate the leak, the LCW pipes (both supply and return) were damaged.  All of the loads to the LCW system have been isolated (RF system).  Bill Rust and Scott from MRI was up there last week to get an idea of how much dirt was in the system and it appears that the piping in the FEL (at least up stairs) does not
contain any dirt.

Tuners have been checked and verified.  We found a bad stepper motor card for positions 1 thru 4 and sent the card back to AES.  They do not have any spares so there is only one stepper card in FL04.

Arc detectors #2 and #7 were found to operate properly.  It was initially reported that #3 had a problem but that was in error (it should have been #2).  The arc heads were bad on both cavities.

Installed missing candy stripe in FL04B14 block 32.  These wires were necessary in order to indicate crowbar status and internal/external interlock status on the CPS.

Continuing to work on Mod Anode Current Monitor system.  One of the problems was that the wrong phone cables were used on the temperature sensor cards.  I am in the process of installing the correct cables.

In the process of troubleshooting the current monitor, I found that the 120v power to the 2.5 amp power supply and the current monitor was not deenergized.  This was a problem due to having no coolant flow and the K5 relay (crydom) should have tripped.  Found the problem to be that the sense for the K5 relay was connected to w/g pressure and not coolant flow.  This problem has been resolved.

In the process of troubleshooting the K5 relay, I went to FL02 to see how it was connected at the bottom of the HPA (since I did not have a correct drawing of K5 at the time).  When I deenergized the 120 volt power to the HPA and the CPS, the waveguide pressure fault energized.  Upon further troubleshooting I found that the W/G pressure interlock is somehow tied to the 120 volt CPS breaker.  This should not be.  I will try to fix this before the week is out.  This problem is only in FL02 (yes, the zone that has been operating).

WBS 8 (I&C):

The procurements for the vacuum controls and ion pump power supplies are ramping up. Thanks to Debbie Griffith for putting the "kits" together for assembly of the power supplies, a total of 40 new power supplies will be fabricated.  At the same time the CEBAF support group (AES) are buying spares, taking advantage of the volume discounts.  The Laser Safety System (LSS) is being "enhanced" to operate with another alignment mode (10 micro sec macro-pulse, 60 Hz repetition rate) and options to run into a hutch or not for each of the user labs. There will be an update of the PLC software to ease alignment of the optical transport system when not in "beam permit". Additionally six monitors were added to the control room to allow continuous monitoring of all of
the user labs. The final version of the surface mount Optical Beam Position Monitor (OBPM) design has been turned into document control for drawings and PC layout. The new LSS access control computer has been installed and is operational, the primary difference is that it can support up to 8 smart card readers.  Software is being prepared for the auto tracking on the optical transport. The schedule is tight but possible to have the collamator, mirror can 2 and mirror can 6 pico motors closed loop on the OBPMs for the Feb. user run.  In the simple case if any of the OBPMs go out of a user defined dead band EPICS will drive the respective pico motor back in to zero position.  Logic may be added depending on performance to follow a give sequence or to limit to moving only one motor at a time.

WBS 9 (Transport):

o The interface document was sent to DULY Research for estimating the effort to design the Injection and Extraction dipoles (GU & GV).  We composed the document to send to Advanced Energy Systems to design the (UV Modified GX), the Reverse Bend (GQ) and the 180 degree bend (GY).
o  We calculated out the new dimensions for a preliminary version of the new Engineering Lattice that David Douglas is working on so that we can place the magnet icons on the layout drawing and get a reality check.

QX (3.125" Quad) Magnets

We are evaluating using a simpler method of construction for yokes that was suggested by a vender as a result of our request for budgetary estimates.  The method is to use ground flat plates and simple pole pieces with an accurately machined pole shape that could be assembled around a fixture that defined the precision dimension of the pole gaps. The parts would then be pinned to one in such a way the gaps would be maintained.  There may be substantial savings in this method.

o  Measurement Probe: The second of the two probe coils, a 100-turn litz coil is complete.  The contract to manufacture the probe body assembly was awarded and the materials brought to the vendor.

o  The Engineering Services Contracts was approved by DOE and established with the vendors.

WBS 10 (Wiggler):

The drawings for the wiggler viewer optical transport are expected to be ready for checkoff by COB Friday.  They should be signed off by Jan. 5.  We received a quote for the beryllium viewers which exceeded our estimate.  We will be looking for other suppliers starting Jan 3.

WBS 11 (Optics):

Our contract at AES on the deformable high reflecting mirror concludes today.  Our work with their staff this week consisted mainly in the areas of document control, and in providing feedback on their latest assembly drawings.

Given the high performance required of our cavity optics, in-house evaluation will be a necessity.  A PO for a laser interferometer was awarded today. We also ordered other inspection instruments and optical quality standards.

The drive laser was serviced this week, and once again delivers specified output.  However, the performance is not as good as we've seen in the past.  This is due to the laser rod being less than the specified diameter (Nd:YLF has a slight solubility in water).  New rods are on order.

We have also been active in writing orders for additional components for the O-BPM system, and have done work to improve the stability of the optical transport system.


No operations  this week. Next operations scheduled will be preops testing at the end of next month in prepartion for the February 2001 user run.

Privacy and Security Notice